# Automatic Stage Lighting Control: Is it a Rule-Driven Process or Generative Task?

### v1

@@ 4.1. Setup
  • Rule-based Method: In reproducing previous methods, we encountered several challenges, as most existing works do not release their code and parameters. Even when attempting to implement them independently, the unreleased predefined light patterns made it impossible to map classification results to their corresponding lighting control outputs. We follow the core idea of Hsiao et al. (Hsiao et al., 2017), which maps music to emotional categories and infers specific lighting patterns based on embeddings in emotional space. For emotion classification, we adopt the method of Aljanaki et al. (Alajanki et al., 2016), training a bidirectional Long Short-Term Memory (LSTM) model (Hochreiter and Schmidhuber, 1997) on their published dataset, DEAM. For light mapping, we utilize the approach of Nijdam (Nijdam, 2009), taking into account both hue and value. We acknowledge that this may introduce some bias for the rule-based methods due to our use of different classification and mapping techniques, as well as datasets. However, we believe we have done our best to achieve a proper reproduction. For lighting control methods based on style classification, we have no means to implement them, as there is currently no established theory guiding how to map style to color.
More details about the model configuration and training process can be found in the appendix. The experiments were conducted using the PyTorch framework (Paszke et al., 1912) on a server running Ubuntu 22.04.5 LTS, equipped with an Intel(R) Xeon(R) Gold 6133 CPU @ 2.50 GHz, two NVIDIA 4090 GPUs, and one NVIDIA A100 GPU.
4.2. Dataset
As mentioned in introduction, there is currently no dataset available for the lighting generation task, which is one reason previous methods have relied on rules. We address this gap by proposing the first stage lighting dataset named ‘RPMC-L2’ (Rock, Punk, Metal, and Core - Livehouse Lighting). This dataset is highly diverse, collected by the authors over the past five years, including contributions from numerous bands, livehouses, and four different capturing devices, covering a wide range of styles such as metalcore, alternative rock, and post-punk. We construct the videos using the proposed method in Section 3.2 from raw footage. Our dataset includes 699 samples, with lengths ranging from 20 seconds to around 5 minutes, making it a relatively medium-sized dataset in MIR. More details can be found in the appendix.
In the experiments, we set the sampling rate to 10 Hz, which means we obtain 10 frames for each second, and our model can process sequences up to 102.4 seconds long. For samples exceeding this length, we randomly select a 102.4-second segment from the entire sequence. We split the dataset into training, validation, and testing sets in an 8:1:1 ratio. To avoid any potential information leakage, we ensure that all shows in the testing set are not included in the training and validation sets. For pre-training, we only use samples from the training and validation sets. In the evaluation, we compare the performance of different models on the testing set.
4.3. Quantitative Analysis
Table 1. Quantitative Results: The best result is indicated in bold, and the second best is indicated by underline. This notation will remain consistent in the following tables.
Methods RMSE MAE
Hue Value Hue Value
Rule-based 48.67 93.84 43.43 86.55
Skip-BART 36.13 60.74 28.71 51.27
from scratch 36.63 68.49 28.83 57.22
w/o skip connection 36.89 68.33 29.44 58.34
w/o light embedding 51.04 67.25 41.50 54.87
In the quantitative experiment, we compare the generation results of different methods against the ground truth. Since our method utilizes random sampling during training, comparing loss and accuracy has limited significance, similar to most generative works (Liang et al., 2024). Therefore, we directly calculate the Root Mean Square Error (RMSE) and Mean Absolute Error (MAE) between the generated results and the ground truth. As hue lies in a cyclic space, we use the distance defined in Eq. 11 for our calculations. The results are presented in Table 1, where it is evident that our method outperforms all ablation studies and the rule-based method in both hue and value, achieving the lowest error. Since the rule-based method does not directly learn from the ground truth, it may introduce some unfairness; we provide further explanations in the human evaluation section.
4.4. Human Evaluation
Due to page limitations, we outline the primary experimental process and results below, while the appendix contains the complete survey questionnaire and the supplementary tables presenting partial analytical results.
4.4.1. Study Setup
To better assess how our results align with human preferences, we designed a questionnaire to collect participants’ feedback on the stage lighting effects created by each object. We recruited 38 respondents through social media platforms and live music venues, with the human evaluation spanning half a month. The participants included 20 males and 18 females aged 18–55 (predominantly 18–30), among whom 9 had professional experience in lighting design, music production, or stage art, and over half were frequent live music attendees. All participants reported normal hearing, normal or corrected-to-normal vision, and no history of color blindness or color weakness.
The questionnaire was structured as follows:
  • 4 objects per group: Ground Truth, Skip-BART, Skip-BART (w/o skip connection), and Rule-based. Here, for the ablation study, we only evaluate Skip-BART without the skip connection. We primarily considered the following factors: first, the efficiency of these components has been widely demonstrated in other fields; second, increasing the number of comparative methods complicates data collection and may discourage user participation, as longer questionnaires can lead to fatigue and confusion when viewing the same lighting video for a particular piece of music. For simplicity, we will refer to Skip-BART (without skip connection) as Ablation Study in this section and the following tables.
  • 6 evaluation dimensions: Emotional Match Between Lighting and Music, Visual Impact, Rhythmic Synchronization Accuracy, Smoothness of Lighting Transitions, Immersive Atmosphere Intensity, and Innovative Surprise (Erdmann et al., 2025).
4.4.2. Data Analysis
4.4.3. Results
As shown in Table 2 and Table 3, Ground Truth receives the highest overall score (M = 4.5, SD = 0.9), closely followed by our Skip-BART(M=4.3, SD=0.9) and the Skip-BART (w/o skip connection)(M=4.1, SD=0.8), both of which are near the GT level. All three methods significantly outperform the rule-based approach(M=2.7, SD=1.3,p¡0.05), a trend that holds for all the metrics. The only metric where Ground Truth does not achieve the best score is in emotion matching, supporting the perspective that the emotional aspects of music may have a limited connection to lighting control (McDonald et al., 2022). We also observe that for the rule-based method, the score of the emotion metrics(M=3.1, SD=1.5) is higher than that of other metrics, which validates our reproduction efforts.
Table 2. Human Evaluation Scores: The six metric scores and overall evaluations of the four objects are expressed as Mean ± Standard Deviation (M ± SD) . The bold text represents the best result, while the underlined text indicates the second-best result.
Method Emotion Impact Rhythm Smoothness Atmosphere Surprise Overall
Ground Truth 4.50±0.93 4.48±0.99 4.61±0.99 4.62±1.07 4.49±0.89 4.34±1.10 4.51±0.88
Skip-BART 4.69±0.87 4.39±0.95 4.50±1.06 4.32±1.12 4.32±0.93 3.83±1.06 4.35±0.87
Ablation Study 4.31±0.94 3.78±0.96 4.54±1.08 4.43±1.12 4.11±0.98 3.50±1.00 4.11±0.84
Rule-based 3.12±1.52 2.65±1.39 2.54±1.47 2.56±1.27 2.77±1.50 2.35±1.40 2.67±1.29
Comparing our Skip-BART with the Skip-BART without skip connection, from which we removed the skip connection, we find that our method consistently outperforms in all metrics except for rhythm($\Delta$M = -0.035, SD=0.12, p=1.00) and smoothness($\Delta$M = -0.11, SD=0.16, p=1.00), with minimal performance gaps. The small gap in rhythm can be attributed to randomness. In terms of smoothness, the skip connection links each frame of music to light directly, which might cause significant changes in light during major transitions in the music. Overall, our method greatly outperforms conventional rule-based methods and achieves performance comparable to that of professional lighting engineers.
Table 3. Pairwise Comparisons of Overall Scores for the Four Objects: The symbols * and ** denote that the significance level (p) for the difference in means ($\Delta$M) is less than 0.05 or 0.01, respectively.
Comparison $\Delta$M SD p
Ground Truth vs. Skip-BART 0.16 0.10 0.724
Ground Truth vs. Ablation Study 0.40 0.10 0.003**
Ground Truth vs. Rule-based 1.84 0.21 ¡ 0.001**
Skip-BART vs. Ablation Study 0.23 0.10 0.152
Skip-BART vs. Rule-based 1.68 0.19 ¡ 0.001**
Ablation Study vs. Rule-based 1.44 0.16 ¡ 0.001**
4.5. Result Visualization
To better present the perceptual quality of the generated lighting sequences, we visualize the output of the four objects in human evaluation, in Fig. 4. Overall, Skip-BART produces sequences that closely resemble the ground truth, indicating that it has learned the underlying variation patterns present in the real data. Notably, the red box highlights a segment transition that Skip-BART successfully identifies, aligning well with the ground truth. Following this transition, both sequences exhibit a brighter light than before. Despite these strengths, Skip-BART still has certain limitations. It occasionally exhibits overly strong local fluctuations and fails to maintain the long-term rhythmic stability observed in the ground truth. This suggests that while the model handles short-term dynamics effectively, its global temporal structure modeling remains an area for improvement. The ablation variant, while visually distinct in color, still demonstrates some ability to follow the underlying temporal structure. In contrast, the rule-based method, relying solely on emotion mapping, tends to produce slow, smooth transitions—typically from blue to cyan to green—which results in outputs that lack the realism and complexity of human-designed lighting.
Figure 4. Visualization of lighting sequences generated by different methods.
5. Discussion and Future Work
Even in the last section, our method demonstrates excellent performance, outperforming conventional rule-based methods and showing significant potential with results closely matching those of real human engineers. However, it remains a relatively preliminary effort. Several areas can be further explored in the future:
  • Leveraging Expert Knowledge: While our end-to-end method demonstrates higher performance and is easier to train, its effectiveness is directly influenced by the quality and quantity of data. With sufficiently large model sizes and data volumes, we can expect scaling laws to enhance the model’s capabilities. Unfortunately, the limited dataset has long been a challenge for the MIR community. Considering future extensions to online real-time control, the model size should not be excessively large. From a deep learning perspective, the pre-training and transfer learning mechanisms employed in this paper are effective strategies for enabling the model to learn from limited data. We can also explore the introduction of external knowledge to support the model; however, we first need to identify the factors that lighting engineers truly consider. Controversial terms like ”emotion” may offer limited assistance. Based on our experience, we find that rhythm, musical sections (such as verses and choruses), and lyrics (focusing on objects and literary imagery rather than merely the theme, as in previous methods) significantly influence lighting design. Consequently, models like beat tracking (Chang and Su, 2024) and melody extraction (section classification) (Zhao, 2025) can be integrated to provide more valuable information. For instance, in our method, experimenting with different hyperparameters of temperature in stochastic temperature-controlled sampling could be beneficial, as choruses may warrant higher randomness than verses. Nevertheless, it is crucial to first study which factors are genuinely related to lighting.
  • Integrating Mature Generation Technologies: This paper introduces a generative model for lighting control, framing it as a creative task. We believe that the successes of generative models can be transferred to this domain. As mentioned, large models have achieved significant success with the aid of scaling laws. Although constructing a large model specifically for the lighting control task may be challenging, developing a general foundation model in music appears feasible. We can directly utilize the trained backbones of some large music models, such as YuE (Yuan et al., 2025) and Suno (Suno, [n. d.]), and also construct foundation models from other fields, such as natural language (Chen et al., 2024). Additionally, other mature technologies, such as Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022), merit exploration.
  • Towards More Practical Scenarios: Similar to most previous studies, our method currently supports only offline primary lighting generation. However, in practice, this is insufficient; we must also consider the features of different venues, which vary in terms of light quantity, position, and configurations (such as color variety and action range). The distribution of colors and light interactions should also be considered, as discussed in an emotion-based approach (Kanno and Fukuhara, 2022), from which we can draw insights. This presents a significant challenge, as we need to design a model capable of simultaneously controlling multiple lights for hue, value, and action. The task becomes even more complex when considering real-time and online control. Training a single model using heterogeneous data from multiple venues poses additional difficulties, though some heterogeneous networks (Wang et al., 2024) may offer potential solutions. Moreover, we encounter cross-domain challenges if the testing venues are not included in the training set, which we believe will require substantial time to address. For the first time, the multiple lighting control task can be considered, meaning that for each frame, we generate all potential lighting instead of relying on a single primary light. To tackle this in an end-to-end manner, we need to extract the hue and value of multiple lights from video. An efficient approach is provided in the appendix to support future research.
6. Conclusion
In this paper, we propose the first end-to-end deep learning method, Skip-BART, for stage lighting generation. We modify the BART model to enable the encoder to extract music features in a single pass and the decoder to generate light hue and intensity auto-regressively, framing ASLC as a generative task. A skip connection mechanism is introduced to enhance the relationship between corresponding music and light frames. Additionally, we incorporate effective deep learning techniques to allow the model to learn from limited data, including transfer learning, pre-training, and restricted stochastic temperature-controlled sampling. To support research in this field, we present a dataset construction method based on raw video and release the first stage lighting dataset, RPMC-L2. Through quantitative analysis and human studies, we demonstrate the efficiency of our method and highlight the limitations of conventional rule-based approaches. Our Skip-BART shows performance closely matching that of real human lighting engineers. The results validate our perspective that music lighting control resembles an art generation task rather than a simple classification and mapping task based on predefined rules. This study aims to provide deeper insights and support for future research in this area.
Appendix
[…]
1
Appendix A Model Configurations
@@
The model configurations in our experiment are shown in Table 4, with the backbone aligned with PianoBART (Liang et al., 2024).
@@
Table 4. Model Configurations: This table provides a comprehensive overview of our Skip-BART settings used in our experiments.
@@
Appendix B Visualization Results
To provide an intuitive understanding of our model’s performance on the stage lighting generation task, we visualize the output of the Skip-BART model alongside the ground-truth lighting control sequences, as shown in Fig. 5a and Fig. 5b. Each example consists of the following three rows:
  • Mel Spectrogram: The Mel spectrogram of the input audio, serving as the model’s conditioning input.
  • Skip-BART: The lighting sequence predicted by the model. The color of each column reflects the model’s prediction of lighting color and brightness.
(a) Example 1
(b) Example 2
Figure 5. Visualization of the stage lighting generation results: The top row displays the input Mel spectrogram, the middle row presents the ground-truth lighting sequence, and the bottom row illustrates the predicted sequence from the Skip-BART model. Each color represents a unique combination of lighting color and brightness over time.
@@ C.2. Audio Processing
  • OpenL3 (Cramer et al., 2019): A deep neural network architecture for extracting high-level auditory embeddings through audio-visual correspondence learning.
@@
Table 5. Audio Features Dimension:$L$ refers to the sequence length of audio.
@@
C.3. Lighting Processing
To facilitate the analysis of lighting conditions, the lighting information is preprocessed in the HSV color space. Specifically, we set the Saturation (S) channel to a constant value [^2] To avoid ambiguity, we use Value or V to refer to the V component in HSV, and value to represent its numerical value here. of 100% to simulate the lighting effect and mitigate environmental influences. The Hue (H) and Value (V) channels are extracted frame-by-frame from the video, with H ranging from [0, 179] (H is cyclic) and V ranging from [0, 255]. For each frame, we compute the distributions (number of pixels for each value) of H and V. To avoid Hue distortion caused by low light intensity, H is extracted only from pixels where V exceeds specified thresholds, set to values in the range {0, 30, 60, 90, 120, 150, 180, 210, 240}. Users can select the threshold values that best suit their specific applications.
C.4. Multiple Hue Extraction
Figure 6. Visualization of a VMM Fitted to Hue Data: The method enables the decomposition of a mixed Hue distribution into several color components.
@@ A VMM models the hue distribution as a weighted sum of $K$ such components:
Appendix D Human Study Details
D.1. Questionnaire Content
@@
      Figure 7. Video Interface Example: The upper-left labels serve as internal markers for video randomization to aid identification, with respondents unaware of their purpose.
@@ D.2. Data Aggregation
Table 6. Median and Plurality Scores: The bold text represents the best result, while the underlined text indicates the second-best result. For the plurality, we retain the highest score when multiple pluralities exist.
Emotion Impact Rhythm Smoothness Atmosphere Surprise Overall
Median
Ground Truth 4.33 4.67 4.67 4.67 4.67 4.50 4.58
Skip-BART 4.67 4.50 4.50 4.33 4.33 4.17 4.42
Ablation Study 4.33 3.83 4.67 4.33 4.17 3.33 4.11
Rule-based 3.17 2.33 2.17 2.50 2.17 1.67 2.33
Plurality
Ground Truth 4.33 5.33 4.67 5.00 5.00 5.00 5.17
Skip-BART 4.67 3.67 5.33 4.33 5.67 4.33 4.72
Ablation Study 4.67 4.00 4.67 4.33 5.00 4.67 4.61
Rule-based 4.00 1.00 1.00 1.00 1.00 1.00 1.00
D.3. Significance Analysis
Table 7. Pairwise Comparisons among Ground Truth (GT), Skip-BART (SB), Ablation Study (AS), and Rule-Based (RB): The symbols * and *** denote that the significance level (p) for the difference in means ($\Delta$M) is less than 0.05 or 0.001, respectively.
Metrics Comparison $\Delta$M SD p Metrics Comparison $\Delta$M SD p
Emotion GT vs SB -0.19 0.15 1.000 Impact GT vs SB 0.09 0.13 1.000
GT vs AS 0.19 0.11 0.598 GT vs AS 0.70 0.15 $<0.001$***
GT vs RB 1.38 0.25 $<0.001$*** GT vs RB 1.83 0.23 $<0.001$***
SB vs AS 0.39 0.14 0.045* SB vs AS 0.61 0.14 0.001***
SB vs RB 1.57 0.23 $<0.001$*** SB vs RB 1.75 0.23 $<0.001$***
AS vs RB 1.18 0.22 $<0.001$*** AS vs RB 1.13 0.18 $<0.001$***
Rhythm GT vs SB 0.11 0.12 1.000 Smoothness GT vs SB 0.30 0.14 0.234
GT vs AS 0.08 0.14 1.000 GT vs AS 0.19 0.16 1.000
GT vs RB 2.07 0.26 $<0.001$*** GT vs RB 2.06 0.23 $<0.001$***
SB vs AS -0.04 0.12 1.000 SB vs AS -0.11 0.16 1.000
SB vs RB 1.96 0.21 $<0.001$*** SB vs RB 1.76 0.23 $<0.001$***
AS vs RB 1.99 0.25 $<0.001$*** AS vs RB 1.87 0.20 $<0.001$***
Atmosphere GT vs SB 0.17 0.14 1.000 Surprise GT vs SB 0.51 0.16 0.014*
GT vs AS 0.38 0.15 0.088 GT vs AS 0.84 0.16 $<0.001$***
GT vs RB 1.72 0.25 $<0.001$*** GT vs RB 1.99 0.23 $<0.001$***
SB vs AS 0.21 0.13 0.619 SB vs AS 0.33 0.17 0.318
SB vs RB 1.55 0.21 $<0.001$*** SB vs RB 1.48 0.21 $<0.001$***
AS vs RB 1.34 0.18 $<0.001$*** AS vs RB 1.15 0.17 $<0.001$***

### v2

@@ Appendix B Detailed Experiment Setup
  • Rule-based Method: In reproducing previous methods, we encountered several challenges, as most existing works do not release their code and parameters. Even when attempting to implement them independently, the unreleased predefined light patterns made it impossible to map classification results to their corresponding lighting control outputs. We follow the core idea of hsiao2017methodology, which maps music to emotional categories and infers specific lighting patterns based on embeddings in emotional space. For emotion classification, we adopt the method of alajanki2016benchmarking, training a bidirectional Long Short-Term Memory (LSTM) model (hochreiter1997long) on their published dataset, DEAM. For light mapping, we utilize the approach of nijdam2009mapping, taking into account both hue and value.
@@
For our Skip-BART model, the detailed configuration in our experiment are shown in Table 4, with the backbone aligned with PianoBART (liang2024pianobart).
@@
Table 4: Model Configurations: This table provides a comprehensive overview of our Skip-BART settings used in our experiments.
@@ C.2 Audio Processing
  • OpenL3 (cramer2019look): A deep neural network architecture for extracting high-level auditory embeddings through audio-visual correspondence learning.
@@
Table 5: Audio Features Dimension:$L$ refers to the sequence length of audio.
@@
C.3 Lighting Processing
To facilitate the analysis of lighting conditions, the lighting information is preprocessed in the HSV color space. The HSV color space separates the brightness (Value) information and the color (Hue) information, which makes the control and generation of lighting more intuitive, independent, and stable. Specifically, we set the Saturation (S) channel to a constant value [^3] To avoid ambiguity, we use Value or V to refer to the V component in HSV, and value to represent its numerical value here. of 255 (100%) to simulate the lighting effect and counteract environmental influences commonly encountered in livehouse settings, ensuring that after factors such as air scattering, fog, or surface reflections reduce color purity, the light maintains vividness and minimizes distortions like fading or whitening. The Hue (H) and Value (V) channels are extracted frame-by-frame from the video, with H ranging from [0, 179] (H is cyclic) and V ranging from [0, 255]. For each frame, we compute the distributions (number of pixels for each value) of H and V. To avoid Hue distortion caused by low light intensity, H is extracted only from pixels where V exceeds specified thresholds, set to values in the range {0, 30, 60, 90, 120, 150, 180, 210, 240}. Users can select the threshold values that best suit their specific applications.
For this study, in order to obtain the audio sequences $s_{a}$, we segment it into overlapping frames $\{x_{1},x_{2},\dots,x_{T}\}$ with a window length of 1 second and a hop size of 0.1 second. To extract the primary light from each image, we first convert each frame into the HSV color space. After flattening, each frame yields a set of $(h,s,v)$ tuples corresponding to its pixels. As described previously, we fix the saturation component $s$, and therefore omit it in subsequent description. We focus on the hue and value components and compute the histogram representations $H_{h,t}\in\mathbb{N}^{180}$ and $H_{v,t}\in\mathbb{N}^{256}$ for frame $t$ as follows:
[…]
where $\mathbf{1}\{\cdot\}$ denotes the indicator function, $h_{t}^{k}\in\{0,\dots,179\}$ and $v_{t}^{k}\in\{0,\dots,255\}$ are the hue and value of pixel $k$ in frame $t$, $K$ is the number of pixels in frame $t$, and $v^{\prime}\in\{0,\dots,255\}$ is a threshold used to exclude low-value pixels.
$y_{t}=[\overline{h_{t}},\overline{v_{t}}]\in\mathbb{Z}^{2}_{*},\quad\text{where}\quad\begin{cases}\overline{h_{t}}=\operatorname{Mode}(H_{h,t})\\ \overline{v_{t}}=\operatorname{Round}(\dfrac{\sum_{n=0}^{255}(n+1)\cdot(H_{v,t})_{n}}{\sum_{n=0}^{255}(H_{v,t})_{n}})\end{cases}$ (12)
Here, we use the mean value to compute the tokenized intensity, and mode to compute the tokenized hue, as hue is a circular variable. This results in a music sequence $x=\{x_{1},x_{2},\ldots,x_{T}\}$ and the corresponding light sequence $y=\{y_{1},y_{2},\ldots,y_{T}\}$. We consider the music sequence $x$ as input and the corresponding light sequence $y$ as output.
C.4 Overall Dataset Settings
As for the overall dataset setting for experiments, we set the sampling rate to 10 Hz, which means we obtain 10 frames for each second, and our model can process sequences up to 102.4 seconds long. For samples exceeding this length, we randomly select a 102.4-second segment from the entire sequence. We split the dataset into training, validation, and testing sets in an 8:1:1 ratio. To avoid any potential information leakage, we ensure that all shows in the testing set are not included in the training and validation sets. For pre-training, we only use samples from the training and validation sets. In the evaluation, we compare the performance of different models on the testing set.
C.5 Multiple Hue Extraction
Figure 4: Visualization of a VMM Fitted to Hue Data: The method enables the decomposition of a mixed Hue distribution into several color components.
@@ A VMM models the hue distribution as a weighted sum of $K$ such components:
Appendix D Visualization Results
To provide an intuitive understanding of our model’s performance on the stage lighting generation task, we visualize the output of the Skip-BART model, ground-truth and other two methods lighting control sequences, as shown in Fig. 5a and Fig. 5b. Each example consists of the following four parts:
  • Skip-BART: The lighting sequence predicted by the Skip-BART model.
  • Ablation Study: The lighting sequence predicted by the BART model.
  • Rule Based: The lighting sequence predicted by rule based method.
For Ground Truth, the color of each column directly represents the actual stage lighting color and brightness at that moment. For other three models, the color of each column reflects the model’s prediction of lighting color and brightness.
(a) Example 1
(b) Example 2
Figure 5: Visualization of lighting sequences generated by different methods. The top row shows the input Mel spectrogram, the middle row is the ground-truth lighting sequence, and the bottom row is the predicted sequence from the Skip-BART model. Each color represents a unique combination of lighting color and brightness over time. The red box in (b) highlights a representative segment where Skip-BART closely matches the temporal lighting structure of the ground truth.
In addition, we present a typical case in Fig. 5b, where the red box highlights a segment transition that Skip-BART successfully identifies, aligning well with the ground truth. Following this transition, both sequences exhibit a brighter light than before. Despite these strengths, Skip-BART still has certain limitations. It occasionally exhibits overly strong local fluctuations and fails to maintain the long-term rhythmic stability observed in the ground truth. This suggests that while the model handles short-term dynamics effectively, its global temporal structure modeling remains an area for improvement. The ablation variant, while visually distinct in color, still demonstrates some ability to follow the underlying temporal structure. In contrast, the rule-based method, relying solely on emotion mapping, tends to produce slow, smooth transitions, typically from blue to cyan to green, which results in outputs that lack the realism and complexity of human-designed lighting.
Overall, Skip-BART is built through transfer learning and pre-training so it learned latent music representations. Also, it is supervised trained, which means it naturally learned the lighting patterns present in the data. Similar to most generative models, it also incorporates controllable randomness through the temperature parameter. This allows the model to move beyond a single deterministic output and introduce variation that may appear some “creativity”.
Appendix E Human Study Details
E.1 Study Setup
We recruited 38 respondents through social media platforms and live music venues, with the human evaluation spanning half a month. The participants included 20 males and 18 females aged 18–55 (predominantly 18–30), among whom 9 had professional experience in lighting design, music production, or stage art, and over half were frequent live music attendees. All participants reported normal hearing, normal or corrected-to-normal vision, and no history of color blindness or color weakness.
The questionnaire was structured as follows:
  • 4 objects per group: Ground Truth, Skip-BART, Skip-BART (w/o skip connection) (i.e., Ablation Study), and Rule-based.
  • 6 evaluation dimensions: Emotional Match Between Lighting and Music, Visual Impact, Rhythmic Synchronization Accuracy, Smoothness of Lighting Transitions, Immersive Atmosphere Intensity, and Innovative Surprise (erdmann2025development).
E.2 Questionnaire Content
@@
      Figure 6: Video Interface Example: The upper-left labels serve as internal markers for video randomization to aid identification, with respondents unaware of their purpose.
@@ E.3 Data Aggregation
Table 6: Median and Plurality Scores: The bold text represents the best result, while the underlined text indicates the second-best result. For the plurality, we retain the highest score when multiple pluralities exist.
Emotion Impact Rhythm Smoothness Atmosphere Surprise Overall Median Ground Truth 4.33 4.67 4.67 4.67 4.67 4.50 4.58 Skip-BART 4.67 4.50 4.50 4.33 4.33 4.17 4.42 Ablation Study 4.33 3.83 4.67 4.33 4.17 3.33 4.11 Rule-based 3.17 2.33 2.17 2.50 2.17 1.67 2.33 Plurality Ground Truth 4.33 5.33 4.67 5.00 5.00 5.00 5.17 Skip-BART 4.67 3.67 5.33 4.33 5.67 4.33 4.72 Ablation Study 4.67 4.00 4.67 4.33 5.00 4.67 4.61 Rule-based 4.00 1.00 1.00 1.00 1.00 1.00 1.00
E.4 Significance Analysis